OpenStudy (anonymous):

Determine the point at which the function r(t) = has a tagent line that is perpendicular to the plane 4x+y-2z=12...

ganeshie8 (ganeshie8):

normal vector of plane is (4,1,-2)

ganeshie8 (ganeshie8):

direction vector of tangent to r(t) is <t^2, 4, 2t> ?

ganeshie8 (ganeshie8):

\[\large \dfrac{t^2}{4} = \dfrac{4}{1} = \dfrac{2t}{-2}\] solve \(\large t\)

OpenStudy (anonymous):

How do we solve for t? By putting in (4,1,-2)? @ganeshie8

ganeshie8 (ganeshie8):

take the last two ratios : \[\large \dfrac{4}{1} = \dfrac{2t}{-2}\] cross multiply and solve \(\large t\)

OpenStudy (anonymous):

t = -4?

ganeshie8 (ganeshie8):

yes! plug that in first two rations and see if it satisfies

OpenStudy (anonymous):

The first two rations? You mean ratios?

ganeshie8 (ganeshie8):

yes

OpenStudy (anonymous):

Okay I did the math and yes it does. :)

OpenStudy (anonymous):

Taking care of neighbors kid..sorry :/